OpenStudy (anonymous):

its a Q from chapter 2 Q) The tangent to the graph of f(x)=2/x at the point (1,2) forms a triangle with the x and y axes. find the perimeter of the triangle, in units. do not simplify your answer.

OpenStudy (anonymous):

dy/dx = -2/x^2 at pt. (1,2) dy/dx = -2 So slope of tangent = -2 Equation of tangent : y - 2 = -2(x-1) which is 2x + y = 4  Length of hypotenuse is sqrt20 So perimeter = 6+ sqrt20 units